# Investigating Test Overfitting on SWE-bench

### v2

@@ 1. Introduction
Our experiments use two LLMs, Claude-3.7 Sonnet and GPT-4o. We use strong existing LLM-based systems as a starting point, obtaining the candidate code $c_{\textrm{new}}$ from Agentless (Xia et al., 2025) and the surrogate test $t_{\textrm{gen}}$ from e-Otter++ (Ahmed et al., 2025). We extend these systems with a test-based code refinement loop that is representative of those found in recent work (Ehrlich et al., 2025; Li et al., 2025c, b, a). The experiments are based on the TDD-bench Verified (Ahmed et al., 2024) benchmark for test-driven development, which is derived from SWE-bench (Jimenez et al., 2024). TDD-bench Verified comprises challenging real-world Python issue resolution tasks. This paper makes the following contributions.
@@ 3. Methodology
    Run an off-the-shelf reproduction test generator (e-Otter++ (Ahmed et al., 2025)) on $d_{\textrm{issue}}$ and $c_{\textrm{old}}$ to obtain an initial candidate test $t_{\textrm{gen}}$.
@@
Since Agentless (Xia et al., 2025) and e-Otter++ (Ahmed et al., 2025) have been described in prior work, here we only briefly discuss characteristics salient to our paper. First, both Agentless and e-Otter++ internally employ some parallel inference scaling. In our pipeline, we only take the single final result selected by each system and feed it into our own test-based code refinement stage, which can be viewed as a form of sequential inference scaling. Second, both Agentless and e-Otter++ internally make use of regression tests $t_{\textrm{old}}\subset c_{\textrm{old}}$ from the original code of the repository. This is a realistic reflection of how a human software engineer might consult regression tests when tackling an issue. At the same time, it only yields limited benefit, since existing tests $t_{\textrm{old}}$ tend not to test the open issue $d_{\textrm{issue}}$.
@@ 4. Results
The experiments in this paper are based on Agentless [^1] https://github.com/OpenAutoCoder/Agentless-generated code patches and e-Otter++ [^2] https://zenodo.org/records/16755224-generated tests. Agentless was evaluated on SWE-bench Verified, and e-Otter++ was evaluated on TDD-Bench Verified, a subset of 449 instances from SWE-bench Verified. We evaluated on all 449 instances. Our primary metric is test overfitting rate. A code patch $c_{\textrm{new}}$ is overfitting tests if it passes $t_{\textrm{new}}$ but fails $t_{\textrm{gold}}$ or $t_{\textrm{old}}$. Dividing the number of instances where that happens by the total instances yields the overfitting rate.
@@ 4.1. Overfitting without Code Refinement (RQ1)
To quantify overfitting without refinement, we run the e-Otter++ generated tests $t_{\textrm{gen}}$ on Agentless code patches $c_{\textrm{new}}$ and see how many pass. Figure 3 shows that with the Claude-3.7-Sonnet model, for 229 instances, the initial code passes the generated tests. Does this mean that the hidden golden tests $t_{\textrm{gold}}$ will pass on these patches? The answer is no. There are 50 instances where $t_{\textrm{gold}}$ failed, rendering them unresolved according to the benchmark. That means the overfitting rate among the 229 instances is $\frac{50}{229}=21.8\%$. For the GPT-4o model, the overfitting rate is 33.0%, even higher than the Claude-3.7-Sonnet model (see Table 1).
@@ 4.2. Overfitting with Code Refinement (RQ2)
Table 2 shows that exposing tests and prompt augmentation resolved 8 samples. Do all these 8 samples help us resolve additional instances? The answer is no. The initial code patch for one instance was already passing on the golden test. And besides these 8, there were 2 instances that went from resolved to non-resolved. So the net gain is +5 instead of +8. That indicates the actual gain from test-based refinement is less than the apparent gain.
@@ 4.3. Impact of Revealing Golden Tests (RQ3)
RQ3 asks what would happen if the test-based code refinement would not use imperfect generated tests $t_{\textrm{gen}}$ but rather golden tests $t_{\textrm{gold}}$. This is a limit study, because in practice, issues do not usually come with oracular executable acceptance tests. Table 3 shows how much overfitting occurs in this setting. Here, the final determination of whether an instance is resolved is based on passing both golden tests $t_{\textrm{gold}}$ and a subset of regression tests $t_{\textrm{old}}$. (While issue-resolution systems such as Agentless have access to the full regression testing suite contained in the original code $c_{\textrm{old}}$, they do not know which subset of those tests are expected to pass after issue resolution.) The overfitting rate, which means passing now-revealed golden tests $t_{\textrm{gold}}$ but failing the relevant subset of regression tests $t_{\textrm{old}}$, is low (13/223=5.8% and 22/194=11.3% with Claude and GPT-4o, respectively). Apart from overfitting, this also shows how well the model performs when golden tests $t_{\textrm{gold}}$ are exposed. When the model passes the golden reproduction tests, the probability that the issue is resolved is high, but not 100%, because regression tests can still fail. The improvement from using golden tests in issue resolution is modest: the number of resolved instances with the Claude model goes up from 229 to 243, around a 3% increase. This experiment emulates a situation where developers follow test-driven development and write tests prior to generating $c_{\textrm{new}}$. Unfortunately, the gains are limited, highlighting that current issue resolution systems often miss the bigger picture.
@@ 5. Discussion
Test-based code refinement gives LLMs a choice to modify both the code (focal function) and/or the test. Table 4 shows that for instances where refinement succeeds, the model mostly elects to modify the focal function instead of the test. For each scenario, the number of test updates is very low compared to focal updates. Also, there are few cases where we see the LLMs try to fix both the focal function and the test. That indicates the model has a preference for focal modification and believes that the tests are mostly perfect.
@@
Figure 4 compares the coverage between unbiased code patches (both $t_{\textrm{gen}}$ and $t_{\textrm{gold}}$ pass) and overfitted patches for Claude-3.7. The median for unbiased patches is 1, while for overfitted patches it is less than 0.8. Also, the mean coverage is higher for unbiased patches (0.95 vs. 0.9). These results indicate that even though both unbiased and overfitted patches pass observed tests $t_{\textrm{gen}}$, overfitted patches have lower coverage.
@@ 7. Conclusion
In this work, we systematically show that test overfitting is still a problem in LLM-based program repair. We also show that test-based refinement may resolve some additional issues but increases overfitting. Therefore, using tests for patch generation needs more deliberation. We share the initial and modified code patches at this link: https://zenodo.org/records/17227319

### v3

@@ Abstract.
journalyear: 2026copyright: ccconference: 34th ACM Joint European Software Engineering Conference and Symposium on the Foundations of Software Engineering; July 05–09, 2026; Montreal, QC, Canadabooktitle: 34th ACM Joint European Software Engineering Conference and Symposium on the Foundations of Software Engineering (FSE Companion ’26), July 05–09, 2026, Montreal, QC, Canadadoi: 10.1145/3803437.3805574isbn: 979-8-4007-2636-1/2026/07ccs: Software and its engineering Automatic programming
@@ 1. Introduction
Our experiments use two LLMs, Claude-3.7 Sonnet and GPT-4o. We use strong existing LLM-based systems as a starting point, obtaining the candidate code $c_{\textrm{new}}$ from Agentless (Xia et al., 2025) and the surrogate test $t_{\textrm{gen}}$ from e-Otter++ (Ahmed et al., 2026). We extend these systems with a test-based code refinement loop that is representative of those found in recent work (Ehrlich et al., 2025; Li et al., 2025c, b, a). The experiments are based on the TDD-bench Verified (Ahmed et al., 2024) benchmark for test-driven development, which is derived from SWE-bench (Jimenez et al., 2024). TDD-bench Verified comprises challenging real-world Python issue resolution tasks. This paper makes the following contributions.
@@ 3. Methodology
    Run an off-the-shelf reproduction test generator (e-Otter++ (Ahmed et al., 2026)) on $d_{\textrm{issue}}$ and $c_{\textrm{old}}$ to obtain an initial candidate test $t_{\textrm{gen}}$.
@@
Since Agentless (Xia et al., 2025) and e-Otter++ (Ahmed et al., 2026) have been described in prior work, here we only briefly discuss characteristics salient to our paper. First, both Agentless and e-Otter++ internally employ some parallel inference scaling. In our pipeline, we only take the single final result selected by each system and feed it into our own test-based code refinement stage, which can be viewed as a form of sequential inference scaling. Second, both Agentless and e-Otter++ internally make use of regression tests $t_{\textrm{old}}\subset c_{\textrm{old}}$ from the original code of the repository. This is a realistic reflection of how a human software engineer might consult regression tests when tackling an issue. At the same time, it only yields limited benefit, since existing tests $t_{\textrm{old}}$ tend not to test the open issue $d_{\textrm{issue}}$.
@@ 4. Results
The experiments in this paper are based on Agentless [^1] https://github.com/OpenAutoCoder/Agentless-generated code patches and e-Otter++ [^2] https://zenodo.org/records/16755224-generated tests. Agentless was evaluated on SWE-bench Verified, and e-Otter++ was evaluated on TDD-Bench Verified, a subset of 449 instances from SWE-bench Verified. We evaluated on all 449 instances. Our primary metric is test overfitting rate. A code patch $c_{\textrm{new}}$ is overfitting tests if it passes $t_{\textrm{gen}}$ but fails $t_{\textrm{gold}}$ or $t_{\textrm{old}}$. Dividing the number of instances where that happens by the total instances yields the overfitting rate.
@@ 4.1. Overfitting without Code Refinement (RQ1)
To quantify overfitting without refinement, we run the e-Otter++ generated tests $t_{\textrm{gen}}$ on Agentless code patches $c_{\textrm{new}}$ and see how many pass. Figure 3 shows that with the Claude-3.7-Sonnet model, for 229 instances, the initial code passes the generated tests. Does this mean that the hidden golden tests $t_{\textrm{gold}}$ will pass on these patches? The answer is no. There are 50 instances where $t_{\textrm{gold}}$ failed, rendering them unresolved according to the benchmark. That means the overfitting rate among the 229 instances is $\frac{50}{229}=21.8\%$. For the GPT-4o model, the overfitting rate is 33.0%, even higher than the Claude-3.7-Sonnet model (see Table 1). Note that a test can overfit due to weaknesses in both the code patches and the tests. However, we did not explicitly differentiate between the two, as doing so would require extensive qualitative studies, which are difficult to scale.
@@ 4.2. Overfitting with Code Refinement (RQ2)
Table 2 shows that exposing tests and prompt augmentation resolved 8 samples. Do all these 8 samples help us resolve additional instances? The answer is no. The initial code patch for one instance was already passing on the golden test. And besides these 8, there were 2 instances that went from resolved to non-resolved. So the net gain is +5 instead of +8. That indicates the actual gain from test-based refinement is less than the apparent gain. The refinement reward’s coverage term may structurally push patches toward fitting $t_{\textrm{gen}}$’s execution path. However, it is also believed that tests with higher coverage are qualitatively better than other tests.
@@ 4.3. Impact of Revealing Golden Tests (RQ3)
RQ3 asks what would happen if the test-based code refinement would not use imperfect generated tests $t_{\textrm{gen}}$ but rather golden tests $t_{\textrm{gold}}$. This is a limit study, because in practice, issues do not usually come with oracular executable acceptance tests. Table 3 shows how much overfitting occurs in this setting. Here, the final determination of whether an instance is resolved is based on passing both golden tests $t_{\textrm{gold}}$ and a subset of regression tests $t_{\textrm{old}}$. (While issue-resolution systems such as Agentless have access to the full regression testing suite contained in the original code $c_{\textrm{old}}$, they do not know which subset of those tests are expected to pass after issue resolution (Chen et al., 2026).) The overfitting rate, which means passing now-revealed golden tests $t_{\textrm{gold}}$ but failing the relevant subset of regression tests $t_{\textrm{old}}$, is low (13/223=5.8% and 22/194=11.3% with Claude and GPT-4o, respectively). Apart from overfitting, this also shows how well the model performs when golden tests $t_{\textrm{gold}}$ are exposed. When the model passes the golden reproduction tests, the probability that the issue is resolved is high, but not 100%, because regression tests can still fail. The improvement from using golden tests in issue resolution is modest: the number of resolved instances with the Claude model goes up from 229 to 243, around a 3% increase. This experiment emulates a situation where developers follow test-driven development and write tests prior to generating $c_{\textrm{new}}$. Unfortunately, the gains are limited, highlighting that current issue resolution systems often miss the bigger picture.
@@ 5. Discussion
Test-based code refinement gives LLMs a choice to modify both the code (focal function) and/or the test. Table 4 shows that for instances where refinement succeeds, the model mostly elects to modify the focal function instead of the test. For each scenario, the number of test updates is low compared to focal updates. Also, there are few cases where the LLMs tries to fix both the focal function and the test. That indicates the model has a preference for focal modification and believes that the tests are mostly perfect.
@@
We compare the coverage between unbiased code patches (both $t_{\textrm{gen}}$ and $t_{\textrm{gold}}$ pass) and overfitted patches for Claude-3.7. The median for unbiased patches is 1, while for overfitted patches it is less than 0.8. Also, the mean coverage is higher for unbiased patches (0.95 vs. 0.9). These results indicate that even though both unbiased and overfitted patches pass observed tests $t_{\textrm{gen}}$, overfitted patches have lower coverage. Execution coverage or static analysis might be used as a warning against overfitting early during patch generation, without relying on hidden golden tests. We leave this for future work.
@@ 7. Conclusion
In this work, we systematically show that test overfitting is still a problem in LLM-based program repair. We also show that test-based refinement may resolve some additional issues but increases overfitting. Therefore, using tests for patch generation needs more deliberation. We share the initial and modified code patches at this link: https://doi.org/10.5281/zenodo.17227318